SQRT. 2x - 3 = X - 3
Answer by stanbon(75887)

SQRT(2x - 3) = X - 3
Square both sides to get:
2x-3 = x^2-6x+9
x^2-8x+12 = 0
(x-6)(x-2)=0
x = 6 or x = 2
-------------------
Checking:
x=6 checks positive
x=2 is extraneous
==========================
Cheers,
Stan H.